# Majority Voting of Doctors Improves Appropriateness of AI Reliance in Pathology

### v1

@@ 1 Introduction
Although J.C.R. Licklider introduced the concept of ‘man-computer symbiosis’ in 1960 (Licklider, 1960), it was not until the last decade that this vision became a more promising reality (Jordan and Mitchell, 2015). By 2023, Artificial Intelligence (AI) has been increasingly discussed to augment humans in critical tasks (Bi et al., 2019; Surden, 2019; Grigorescu et al., 2020). Especially in the medical domain of pathology, AI has been showcased to increase doctors’ accuracy and speed (Litjens et al., 2016; Lindvall et al., 2021; Van der Laak et al., 2021; Ba et al., 2022), consistency (Balkenhol et al., 2019; Van Bergeijk et al., 2023), and confidence (Gu et al., 2023b). However, because pathology AI was often trained from a limited dataset, its performance varied while being applied to data from new patients and hospitals (Stacke et al., 2021; Aubreville et al., 2021; Gu et al., 2021). As such, it is critical for pathologists to develop appropriate reliance while collaborating with AI, i.e., to appropriately accept correct AI recommendations and reject the wrong ones.
@@
This research aims to provide the validation of the majority voting on enabling the appropriate AI reliance in pathology decision-making, with a focus on a visual search task of detecting “mitosis,” a critical histology pattern for tumor grading (Collan et al., 1996; Meyer et al., 2005). 32 pathologists from ten institutions participated in a multi-stage user study, where they detected mitoses manually, first, and with AI assistance after a wash-out period. Here, the majority voting decisions were synthesized according to the AI-assisted decisions from an odd number of randomly-selected pathologist participants. Two metrics were employed to measure the appropriateness of AI reliance: “relative AI reliance” and “relative self-reliance” (Schemmer et al., 2023). The result showed that the majority voting decisions from as few as three pathologists showed significantly higher relative AI reliance ($\sim 9\%$ increase) and relative self-reliance ($\sim 31\%$ increase), compared to one pathologist collaborating with AI, respectively. The precision and recall of majority voting decisions also increased: Those from three AI-assisted pathologists could achieve a mean precision of 0.902 and a recall of 0.843. As a comparison, the mean precision and recall for one-pathologist-AI collaboration were 0.824 and 0.817, respectively. Furthermore, the majority voting decisions could also have a higher chance of achieving complementary team performance in the recall.
@@ 1.1 Contributions
This research showcases that majority voting can enable appropriate AI reliance for pathology decision-making. Throughout a multi-institutional study amongst 32 pathology professionals, this research presents the effectiveness of majority voting in a high-stakes medical task, which can ultimately benefit patient management. This signifies a transformation from the traditional one-human-AI collaboration to harnessing group decision-makings of AI-assisted medical professionals. While our primary focus has been on pathology, we envision that the insights of this study can be generalized to other high-stakes visual search tasks, such as detecting explosives from X-ray scans or disaster assessment from satellite imagery.
@@ 2.1 Enabling Appropriate AI Reliance
Mainstream research focuses on strategies to reduce over-reliance. For instance, the cognitive forcing function, which prompts users to think analytically before decision-making, has shown promise (Buçinca et al., 2021). Similarly, altering the interaction speed, where enlonging the AI response time, can instigate users’ reflective thinking. Therefore, over-reliance incidents could be reduced (Park et al., 2019; Rastogi et al., 2022; Lebedeva et al., 2023). Other approaches aim to enhance users’ onboarding process, such as improving AI literacy (Lai et al., 2020; Long and Magerko, 2020; Leichtmann et al., 2023), where users are informed of AI details (Cai et al., 2019b; Jacobs et al., 2021a, b). However, translating these approaches to the medical domain may encounter two challenges. Firstly, introducing cognitive forcing functions or altering interaction speed could develop ‘algorithm aversion,’ especially when medical tasks are time-sensitive (Efendić et al., 2020; Fogliato et al., 2022). Secondly, the efficacy of enhancing AI literacy also appeared marginal, possibly because of the difficulties in educating users within a limited timeframe (Lai et al., 2020; Leichtmann et al., 2023).
Besides these, another popular approach is XAI, aiming to reduce over-reliance by enabling users to understand AI’s reasoning (Bussone et al., 2015; Lai et al., 2020; Zhang et al., 2020; Bansal et al., 2021). Nonetheless, numerous studies have failed to observe the anticipated effectiveness of XAI (Schemmer et al., 2022): The potential benefits of XAI may be offset by the cognitive efforts of interpreting them (Vasconcelos et al., 2023). Given the already high cognitive demands of medical professionals, this might result in XAI being less referred to, countering its potential benefits.
@@ 2.2 Decision-Making Processes by Multiple Medical Professionals
Deriving medical decisions has been introduced to and standardized in the healthcare domain since the 1950s (Murphy et al., 1998; Black et al., 1999). Traditionally, there are three primary approaches: (1) the Delphi method, (2) the nominal group technique, and (3) the consensus development conference. These methods have been historically utilized in medical decision-making and guideline formulation (Murphy et al., 1998; Black et al., 1999).
@@
Regarding employing multiple pathologists to make decisions with AI assistance, recent research has implemented majority voting among three doctors to label data for AI development (Bertram et al., 2019; Aubreville et al., 2020). This process involves two pathologists independently annotating data with AI assistance. If there were conflicts, a third pathologist joined and annotated again to formulate majority. This approach resembles the Delphi method, where two group members first formed their judgment independently, and the conflicts were addressed by a third pathologist’s input. However, these studies primarily focus on data labeling for AI development, often featuring non-diverse participant pools (typically three pathologists) with similar backgrounds. Furthermore, they lack analyses of AI reliance metrics, which is critical for assessing the method’s quality.
Our study aims to fill the gap by providing a comprehensive experiment and evaluation of the majority voting on pathology decision-making. To achieve this, we hosted a multi-stage, multi-institutional user study involving 32 pathologists with varied experience levels. We examined the quality of these majority-voted, AI-assisted decisions from two angles: reliance on AI and correctness. Additionally, we evaluated the potential costs of employing this method, providing empirical data for future research in HCI, cognitive science, and medicine.
@@ 2.3 Human-AI Collaboration in Pathology
AI, particularly deep learning, holds promise in performing pathology tasks, ranging from conducting high-level diagnoses (e.g., prostate cancer grading (Pantanowitz et al., 2020)), to detecting low-level pathological patterns (e.g., cell detection (Amgad et al., 2022)). Notably, recent studies have suggested that AI’s performance is on par with human experts in specific pathology tasks (Hekler et al., 2019; Zhang et al., 2019; Wang et al., 2023). Due to legislative and ethical concerns, AI algorithms and software cannot replace pathologists’ examinations (Chauhan and Gullapalli, 2021; Veale and Zuiderveen Borgesius, 2021). Instead, they are regarded as medical devices to assist doctors [^1] https://www.fda.gov/media/145022/download, with one designed for prostate cancer pathology receiving the first official approval in 2021222https://www.fda.gov/news-events/press-announcements/fda-authorizes-software-can-help-identify-prostate-cancer. By 2023, a plethora of human-AI collaborative tools have been introduced, claiming improvements in speed and correctness in detecting pathological patterns (Litjens et al., 2016; Lindvall et al., 2021; Van der Laak et al., 2021; Ba et al., 2022), inter-observer consistencies (Balkenhol et al., 2019; Van Bergeijk et al., 2023), mental workload and confidence (Gu et al., 2023b, c), compared to pathologists examining manually.
Among these improvements, of particular interest is achieving complementary team performance, where pathologist-AI collaboration could outperform both the pathologist and AI (Bansal et al., 2021): In 2016, Wang et al.  reported that combining AI and pathologist’s predictions could reduce error rates in breast cancer classification, theoretically confirming such complementary team performance (Wang et al., 2016). Despite this theoretical backing, there is a lack of empirical evidence to support it for the pathology domain. In the general domain, several studies have failed to observe the task accuracy improvement in human-AI collaboration compared to AI alone (Bansal et al., 2021; Lai and Tan, 2019; Jacobs et al., 2021b). This issue may stem from users’ accepting incorrect AI recommendations, a factor that can significantly impact the outcome of human-AI collaboration (Kaur et al., 2020; Cao and Huang, 2022; Vasconcelos et al., 2023).
@@ 3.1 Task Selection & Generalizability of the Task
According to the 2021 World Health Organization central nervous system tumor classification guidelines, mitosis serves as a critical diagnostic criterion for grading numerous brain tumors, such as IDH-mutant astrocytoma, oligodendroglioma, and ependymoma (Louis et al., 2021). Going beyond pathology, the task selected involves assisting domain professionals in visually detecting small-dimensional, sparsely distributed patterns, which can benefit from AI augmentation. Some of these tasks are time-sensitive but have innate high stakes, such as scanning explosives at security checkpoints by X-ray imaging (Wolfe et al., 2007), or disaster assessment from satellite imagery (Morrison et al., 2023).
@@ 4.1 Participants
Participants were recruited through sending emails to the mailing list and snowball recruitment. As a result, 32 pathology professionals from 10 medical centers in the United States took part in both study stages, including 12 specialist pathologists, six general pathologists, ten residents, and four medical students [^3] The four medical student participants underwent a 45-minute training session overseen by a specialist pathologist before participating, to ensure their familiarity with the mitosis detection task.. The demographic information of participants is shown in A.
@@ 4.2 Study Procedure
    Random image transforms: Including random flipping (vertical and/or horizontal) and random rotation (randomly chosen from {${}^{\circ}$, ${}^{\circ}$, ${}^{\circ}$, and ${}^{\circ}$}). For instance, the image shown in Figure 3(b) was rotated ${}^{\circ}$ anti-clockwise from that in Figure 3(a).
@@
Figure 3: Screenshots of the mitosis study websites: (a) The manual mitosis detection website in the stage 1 study. The user could left-click on the image to leave a mark for each mitosis detected (1⃝ – 3⃝). (b) The AI-assisted mitosis detection website in the stage 2 study. The interface added 1⃝ the AI recommendation box; 2⃝ “Show AI” switch, where the user could toggle on/off AI recommendations; 3⃝ “AI Sensitivity” slider, where the user could adjust the sensitivity of AI based on their preference; 4⃝ a warning message to remind users not relying on AI. (c) The website in stage 2 also provided an XAI evidence card for each AI recommendation. Each XAI evidence card included 1⃝ a saliency map; 2⃝ confidence level, including a probability score and a trust score; 3⃝ a bar plot for subclass probability; and 4⃝ similar examples. (d) After the user finishes examining all images, an evaluation page will inform the performance metrics to the participant.
@@ 4.3.1 Stage 1: Manual Mitosis Detection
This interface only showed participants the images and logged their interactions (Figure 3(a)). If the user found a mitosis, they could left-click on where it resided to leave a mark (Figure 3(a) 1⃝ – 3⃝). The user could go to the next image after examining one. However, they could not return to the previous image to ensure a precise measurement for time consumption. After all images were examined, a status page (Figure 3(d)) was displayed to inform the participant of the performance of their mitosis detection.
@@ 4.3.2 Stage 2: AI-Assisted Mitosis Detection
The AI model used in this stage was an EfficientNet-b3 Convolutional Neural Network (CNN), trained from a meningioma mitosis dataset (Tan and Le, 2019; Gu et al., 2023a). The website displayed AI mitosis detections through recommendation boxes (Figure 3(b)). Additionally, the following components were included to mitigate the negative reduce of improper AI reliance:
@@
    Warning messages: A “black-box” style444… the highest safety-related warning assigned by the U.S. Food and Drug Administration (Delong and Preuss, 2019). warning message was presented in the tutorial video, suggesting that the users should always rely on their judgments. The message was also shown in a highlighted box on the website (Figure 3(b) 4⃝).
@@
    XAI: Each AI recommendation was accompanied by an XAI evidence card. The user could right-click on the AI recommendation box to see the XAI evidence card on-demand. Four popular XAI techniques were included following previous work (Evans et al., 2022), including:
@@
      Confidence level: Including a probability score and a trust score (Zhang et al., 2020). The trust score was the geometric mean of noise (Ayhan and Berens, 2022) and random AI variances (Gal and Ghahramani, 2016) of the AI prediction.
@@
    Counterfactual explanations were not used because of the low quality of the retrieval results.
@@
    Personalized AI adjustments: The user could toggle on/off AI recommendations by interacting with the “Show AI” switch (Figure 3(b) 2⃝) (i.e., AI on-request, suggested by Gaube et al. (2021)) and adjust the AI sensitivity (Figure 3(b) 3⃝) according to their preferences. The website provided five AI sensitivity settings for users: “lowest,” ‘low,” “medium,” “high,” and “highest.” A higher sensitivity would include more AI recommendations with lower probabilities.

### v3

@@ Abstract
{highlights}
Multi-institutional user study with medical doctors from diverse backgrounds
Two quantifiable metrics for evaluating appropriateness of AI reliance
Majority voting improves appropriateness of AI reliance in pathology decisions
Increased appropriateness elevates precision and recall
Grouped pathologists improve qualities in visual decision-making
@@ 1 Introduction
Although J.C.R. Licklider introduced the concept of ‘man-computer symbiosis’ in 1960 (Licklider, 1960), it was not until the last decade that this vision became a more promising reality (Jordan and Mitchell, 2015). By 2023, Artificial Intelligence (AI) has been increasingly discussed to augment humans in critical tasks (Bi et al., 2019; Surden, 2019; Grigorescu et al., 2020). Especially in the medical domain of pathology, AI has been showcased to increase doctors’ accuracy and speed (Litjens et al., 2016; Lindvall et al., 2021; Van der Laak et al., 2021; Ba et al., 2022), consistency (Balkenhol et al., 2019; Van Bergeijk et al., 2023), and confidence (Gu et al., 2023b). However, because pathology AI was often trained from a limited dataset its performance varied while being applied to data from new patients and hospitals (Stacke et al., 2020; Aubreville et al., 2021; Gu et al., 2021). As such, it is critical for pathologists to develop appropriate reliance while collaborating with AI, i.e., to appropriately accept correct AI recommendations and reject the wrong ones.
@@
This research aims to provide the validation of the majority voting on enabling the appropriate AI reliance in pathology decision-making, with a focus on a visual search task of detecting “mitosis,” a critical histology pattern for tumor grading (Collan et al., 1996; Meyer et al., 2005). 32 medical professionals in pathology from ten institutions participated in a multi-stage user study, where they detected mitoses manually, first, and with AI assistance after a wash-out period. Here, the majority voting decisions were synthesized according to the AI-assisted decisions from an odd number of randomly-selected pathologist participants. Two metrics were employed to measure the appropriateness of AI reliance: “relative AI reliance” and “relative self-reliance” (Schemmer et al., 2023). The result showed that the majority voting decisions from as few as three pathologists showed significantly higher relative AI reliance ($\sim 9\%$ increase) and relative self-reliance ($\sim 31\%$ increase), compared to one pathologist collaborating with AI, respectively. The precision and recall of majority voting decisions also increased: Those from three AI-assisted pathologists could achieve a mean precision of 0.902 and a recall of 0.843. As a comparison, the mean precision and recall for one-pathologist-AI collaboration were 0.824 and 0.817, respectively. Furthermore, the majority voting decisions could also have a higher chance of achieving super-AI performance in the recall.
@@ 1.1 Contributions
This research showcases that majority voting can enable appropriate AI reliance for pathology decision-making. Throughout a multi-institutional study amongst 32 pathology professionals, this research presents the effectiveness of majority voting in a high-stakes medical task, which can ultimately benefit patient management. This signifies a transformation from the traditional one-human-AI collaboration to harnessing group decision-makings of AI-assisted medical professionals. While our primary focus has been on pathology, we envision that the insights of this study can have broader implications for leveraging collective human-AI decision-making in other high-stakes visual search tasks, such as detecting explosives from X-ray scans or disaster assessment from satellite imagery for emergency response efforts.
@@ 2.1 Enabling Appropriate AI Reliance
In the context of enabling appropriate AI reliance, this is a tendency in research to study mechanisms and counter-measures for over-reliance. For instance, the cognitive forcing function, which prompts users to think analytically before decision-making, has shown promise (Buçinca et al., 2021). Similarly, altering the interaction speed, where enlonging the AI response time, can instigate users’ reflective thinking. Therefore, over-reliance incidents could be reduced (Park et al., 2019; Rastogi et al., 2022; Lebedeva et al., 2023). Other approaches aim to enhance users’ onboarding process, such as improving AI literacy (Lai et al., 2020; Long and Magerko, 2020; Leichtmann et al., 2023), where users are informed of AI details (Cai et al., 2019b; Jacobs et al., 2021a, b). However, translating these approaches to the medical domain may encounter two challenges. Firstly, introducing cognitive forcing functions or altering interaction speed could develop ‘algorithm aversion,’ especially when medical tasks are time-sensitive (Efendić et al., 2020; Fogliato et al., 2022). Secondly, the efficacy of enhancing AI literacy also appeared marginal, possibly because of the difficulties in educating users within a limited timeframe (Lai et al., 2020; Leichtmann et al., 2023).
Besides these, another popular approach is XAI, aiming to reduce over-reliance by enabling users to understand AI’s reasoning (Bussone et al., 2015; Lai et al., 2020; Zhang et al., 2020; Bansal et al., 2021). Nonetheless, numerous studies have failed to observe the anticipated effectiveness of XAI (Schemmer et al., 2022): The potential benefits of XAI may be offset by the cognitive efforts of interpreting them (Vasconcelos et al., 2023). Given the already high cognitive demands of medical professionals, this might result in XAI being less referred to, countering its potential benefits. This issue of appropriateness usage of XAI in medicine was raised by Holzinger et al. (2019). Further research suggested causability, an ability of an explanation that can enable casual understanding of medical experts, should also considered and measured to achieve better efficiency, effectiveness, and user satisfaction (Holzinger et al., 2020; Plass et al., 2023).
@@ 2.2 Decision-Making Processes by Multiple Medical Professionals
Different from one medical professional examining the specimens (Pohn et al., 2019; Pena and Andrade-Filho, 2009), decision-making processes by medical professionals require communication, discussion, and result sharing (Murphy et al., 1998; Black et al., 1999). Nowadays, there are three primary approaches: (1) the Delphi method, (2) the nominal group technique, and (3) the consensus development conference. These methods have been historically utilized in medical decision-making and guideline formulation (Murphy et al., 1998; Black et al., 1999).
@@
Regarding employing multiple pathologists to make decisions with AI assistance, recent research has implemented majority voting among three doctors to label data for AI development (Bertram et al., 2019; Aubreville et al., 2020). This process involves two pathologists independently annotating data with AI assistance. If there were conflicts, a third pathologist joined and annotated again to formulate majority (Montezuma et al., 2023). However, these studies primarily focus on data labeling for AI development, often featuring non-diverse participant pools (typically three pathologists) with similar backgrounds. Furthermore, they lack analyses of AI reliance metrics, which is critical for assessing the method’s quality.
Our study aims to fill the gap by providing a comprehensive experiment and evaluation of the majority voting on pathology decision-making. To achieve this, we hosted a multi-stage, multi-institutional user study involving 32 medical professionals in pathology with varied experience levels. We examined the quality of these majority-voted, AI-assisted decisions from two angles: reliance on AI and correctness. Additionally, we evaluated the potential costs of employing this method, providing empirical data for future research in HCI, cognitive science, and medicine.
@@ 2.3 Human-AI Collaboration in Pathology
AI, particularly deep learning, holds promise in performing a wide variety of pathology tasks to assist medical professionals (Regitnig et al., 2020), ranging from conducting high-level diagnoses (e.g., prostate cancer grading (Pantanowitz et al., 2020)), to detecting low-level pathological patterns (e.g., cell detection (Amgad et al., 2022)). Notably, recent studies have suggested that AI’s performance is on par with human experts in specific pathology tasks (Hekler et al., 2019; Zhang et al., 2019; Wang et al., 2023). Due to legislative and ethical concerns, AI algorithms and software cannot replace pathologists’ examinations (Chauhan and Gullapalli, 2021; Veale and Zuiderveen Borgesius, 2021). Instead, they are regarded as medical devices to assist doctors [^1] https://www.fda.gov/media/145022/download, with one designed for prostate cancer pathology receiving the first official approval in 2021222https://www.fda.gov/news-events/press-announcements/fda-authorizes-software-can-help-identify-prostate-cancer. By 2023, a plethora of human-AI collaborative tools have been introduced, demostrating improvements in speed and correctness in detecting pathological patterns (Litjens et al., 2016; Lindvall et al., 2021; Van der Laak et al., 2021; Ba et al., 2022), inter-observer consistencies (Balkenhol et al., 2019; Van Bergeijk et al., 2023), mental workload and confidence (Gu et al., 2023b, c), compared to pathologists examining manually.
Among these improvements, of particular interest is achieving complementary team performance, where pathologist-AI collaboration could outperform both the pathologist and AI (Bansal et al., 2021): In 2016, Wang et al.  reported that combining AI and pathologist’s predictions could reduce error rates in breast cancer classification, theoretically confirming the existence of such complementary team performance (Wang et al., 2016). Despite this theoretical backing, there is a lack of empirical evidence to support it for the pathology domain. In the general domain, several studies have failed to observe the task accuracy improvement in human-AI collaboration compared to AI alone (Bansal et al., 2021; Lai and Tan, 2019; Jacobs et al., 2021b). This issue may stem from users’ accepting incorrect AI recommendations, a factor that can significantly impact the outcome of human-AI collaboration (Kaur et al., 2020; Cao and Huang, 2022; Vasconcelos et al., 2023).
@@ 3.1 Task Selection & Generalizability of the Task
According to the 2021 World Health Organization central nervous system tumor classification guidelines, mitosis serves as a critical diagnostic criterion for grading numerous brain tumors, such as IDH-mutant astrocytoma, oligodendroglioma, and ependymoma (Louis et al., 2021). Going beyond mitoses, pathologists may also be required to detect small-scale, sparsely distributed patterns in large scans, such as finding small tumor deposits within lymph nodes in breast cancer or malignant melanoma (Regitnig et al., 2020). In a more general context, similar visual search tasks also exist in high-stakes domains where AI assistance could be valuable. For instance, security personnel must swiftly identify potential threats like explosives in X-ray scans (Wolfe et al., 2007), and emergency responders rely on timely assessments of disaster impacts from satellite imagery (Morrison et al., 2023).
@@ 4.1 Participants
Participants were recruited through sending emails to the mailing list and snowball recruitment. As a result, 32 pathology professionals from 10 medical centers in the United States took part in both study stages, including 12 specialist pathologists, six general pathologists, ten residents, and four medical students [^3] The four medical student participants underwent a 45-minute training session overseen by a specialist pathologist before participating, to ensure their familiarity with the mitosis detection task.. The demographic information of participants is shown in the supplemental material.
@@ 4.2 Study Procedure
    Random image transforms: Including random flipping (vertical and/or horizontal) and random rotation (randomly chosen from {0, 90, 180, and 270}). For instance, the image shown in Figure 3(b) was rotated 270 anti-clockwise from that in Figure 3(a).
@@
Figure 3: Screenshots of the mitosis study websites: (a) The manual mitosis detection website in the stage 1 study. The user could left-click on the image to leave a mark for each mitosis detected (① – ③). (b) The AI-assisted mitosis detection website in the stage 2 study. The interface added ① the AI recommendation box; ② “Show AI” switch, where the user could toggle on/off AI recommendations; ③ “AI Sensitivity” slider, where the user could adjust the sensitivity of AI based on their preference; ④ a warning message to remind users not relying on AI. (c) The website in stage 2 also provided an XAI evidence card for each AI recommendation. Each XAI evidence card included ① a saliency map; ② confidence level, including a probability score and a trust score; ③ a bar plot for subclass probability; and ④ similar examples. (d) After the user finishes examining all images, an evaluation page will inform the performance metrics to the participant.
@@ 4.3.1 Stage 1: Manual Mitosis Detection
This interface only showed participants the images and logged their interactions (Figure 3(a)). If the user found a mitosis, they could left-click on where it resided to leave a mark (Figure 3(a) ① – ③). The user could go to the next image after examining one. However, they could not return to the previous image to ensure a precise measurement for time consumption. After all images were examined, a status page (Figure 3(d)) was displayed to inform the participant of the performance of their mitosis detection.
@@ 4.3.2 Stage 2: AI-Assisted Mitosis Detection
The AI model used in this stage was an EfficientNet-b3 Convolutional Neural Network (CNN), trained from a meningioma mitosis dataset (Tan and Le, 2019; Gu et al., 2023a). The website displayed AI mitosis detections through recommendation boxes (Figure 3(b)). Additionally, following previous works, we included four components to mitigate the negative influence of improper AI reliance:
@@
    Warning messages: A “black-box” style444… the highest safety-related warning assigned by the U.S. Food and Drug Administration (Delong and Preuss, 2019). warning message was presented in the tutorial video, suggesting that the users should always rely on their judgments. The message was also shown in a highlighted box on the website (Figure 3(b) ④).
@@
    XAI: Each AI recommendation was accompanied by an evidence card which attempts to provide XAI assistance (Plass et al., 2023). The user could right-click on the AI recommendation box to see the XAI evidence card on-demand. Four popular XAI techniques were included following previous work (Evans et al., 2022), including:
@@
      Confidence level: Including a probability score and a trust score (Zhang et al., 2020). The trust score was the geometric mean of noise (Ayhan and Berens, 2018) and random AI variances (Gal and Ghahramani, 2016) of the AI prediction.
@@
    Counterfactual explanations were not used because of the low quality of the retrieval results achieved by the our AI model.
@@
    Personalized AI adjustments: The user could toggle on/off AI recommendations by interacting with the “Show AI” switch (Figure 3(b) ②) (i.e., AI on-request, suggested by Gaube et al. (2021)) and adjust the AI sensitivity (Figure 3(b) ③) according to their preferences. The website provided five AI sensitivity settings for users: “lowest,” ‘low,” “medium,” “high,” and “highest.” A higher sensitivity would include more AI recommendations with lower probabilities.